And unknown sample is being evaluated in lab. What is the specific heat capacity of the compound if it requires 313.79 J to raise the temperature of 89.08 grams of the unknown from 15.2 °C to 50.5 °C.
Answer #1

Specific heat capacity is basically the amount of heat required to raise the temperature of 1 kg of given substance by 1K.

Formula for specific heat capacity is ,

Q = m*c*T

Where, Q is the amount of heat required in Joules

m is the mass of substance

c is the specific heat capacity

T is the change in temperature

Now, it is given that

Heat required Q = 313.79 J

m = 89.08 g = 89.08/1000 Kg = 0.08908 Kg

T = 50.5 - 15.2 = 35.3°C = 35.3 K ( because for change units can be put directly it will cancel factor of 273 )

Substituting values in above formula,

313.79 = 0.08908 * c * 35.3

c = 313.79 / ( 0.08908 * 35.3 ) = 99.789 J/ Kg Kelvin ( answer )

I hope you will find the answer useful, for any further clarification please ask in comments.